# GTPO and GRPO-S: Token and Sequence-Level Reward Shaping with Policy Entropy

### v5

@@ Abstract
Reinforcement learning (RL) is a pivotal task for enhancing Large Language Model (LLM) reasoning. Conventional algorithms, however, typically adhere to a coarse-grained credit assignment paradigm, applying a uniform reward to all tokens in a sequence—a critical flaw in long-chain reasoning tasks. In this paper, we address this challenge and propose Dynamic Entropy Weighting, a novel mechanism that facilitates fine-grained rewards through two new algorithms: Group Token Policy Optimization (GTPO), which assigns an entropy-weighted reward to each token, and the analogous algorithm Sequence-Level GRPO (GRPO-S). Our approach is founded on the hypothesis that high policy entropy within a reasoning path is a powerful heuristic for ”cognitive effort” at pivotal junctures, which can be repurposed into a learning signal. By repurposing policy entropy for reward shaping, we achieve true per-token credit assignment. Experimental results across challenging reasoning benchmarks validate the superiority of our approach, showing our methods significantly outperform a strong DAPO baseline and confirming our entropy-weighting mechanism as the key driver of this performance boost.
@@ 1 Introduction
The reasoning capabilities of Large Language Models (LLMs) have evolved profoundly, transitioning from pattern recognition to simulating deeper cognitive processes (Guo et al., 2025; Wei et al., 2022; Zhang et al., 2022; Achiam et al., 2023). This leap is evidenced by state-of-the-art performance on formidable tasks like advanced mathematics and competitive coding, where models learn complex, process-oriented behaviors such as self-verification and iterative refinement (Chen et al., 2025; Lewkowycz et al., 2022; Li et al., 2022; Madaan et al., 2023). At the heart of this paradigm shift lies large-scale Reinforcement Learning (RL), the core technology enabling models to acquire dynamic, multi-step problem-solving strategies beyond the scope of static supervised learning (Zhang et al., 2025b; Wang et al., 2016; Schulman et al., 2017; Ouyang et al., 2022).
This evolution in alignment algorithms has followed a clear trajectory: a persistent quest for greater simplicity and efficiency (Mnih et al., 2015; Amini et al., 2024). Early Reinforcement Learning from Human Feedback (RLHF) pipelines, often built on Proximal Policy Optimization (PPO), were powerful but notoriously cumbersome, necessitating separate reward and value models that introduced computational overhead and training instabilities (Schulman et al., 2017; Christiano et al., 2017; OpenAI Spinning Up, 2018; Gao et al., 2023). In response, the field gravitated towards more direct paradigms. Direct Preference Optimization (DPO) marked a milestone by bypassing explicit reward modeling (Rafailov et al., 2023; 2024; Lai et al., 2024). This trend culminated in value-function-free methods like Group Relative Policy Optimization (GRPO), which simplifies the optimization landscape by using a group’s average reward as an advantage baseline, making it ideal for fine-tuning massive models ((Shao et al., 2024; Yu et al., 2025); see Appendix A for a detailed review of alignment algorithm evolution).
However, this pursuit of efficiency has created a critical bottleneck: coarse-grained credit assignment (Wei et al., 2022). Algorithms like GRPO assign a uniform reward to every token in a sequence based solely on the final outcome (Shao et al., 2024). This limitation becomes profound in tasks requiring long-chain reasoning. For instance, a sequence with dozens of correct logical steps may receive zero reward for a single final error, penalizing correct and incorrect reasoning alike (Yang et al., 2025; Wang et al., 2025). Conversely, a sequence that reaches a correct answer through flawed or guessed intermediate steps is fully rewarded (Sutton and Barto, 2018; Zhang et al., 2025a). This sparse, imprecise feedback constrains learning, creating an efficiency-precision trade-off where imprecise credit assignment is the primary barrier to progress (Bansal et al., 2023).
@@
Figure 1: Conceptual illustration of reward assignment. (a) Traditional methods assign a uniform reward based on the final outcome. In contrast, our methods use Dynamic Entropy Weighting to refine credit assignment: (c) GTPO rewards high-entropy tokens in correct sequences while suppressing them in incorrect ones, and (d) GRPO-S rewards correct sequences with higher average entropy while penalizing incorrect paths. Yielding superior performance (b).
@@
To distinguish the quality of steps in answers and address the coarse-grained credit assignment bottleneck, this paper uses policy entropy to increase the rewards for key reasoning steps in correct answers and reduce the suppression of effective exploration signals in incorrect answers. The central hypothesis is that moments of high policy entropy within a reasoning sequence are not random noise but strong correlates of pivotal reasoning junctures (Cheng et al., 2025; Cui et al., 2025). When a model selects between multiple valid mathematical theorems or constructs a complex logical connective, its uncertainty—as measured by entropy—naturally increases. This policy entropy, traditionally viewed as a measure of model indecision, can be repurposed as a powerful heuristic for cognitive effort (Haarnoja et al., 2018; Lindsay, 2020). In successful paths, it signals a moment of valuable exploration to be reinforced; in unsuccessful paths, it can help the policy break from incorrect thinking. This principle motivates the proposal of Dynamic Entropy Weighting, a novel framework that reshapes the reward signal to be proportional to token-level or sequence-level entropy, thereby focusing the policy gradient on the most critical decision points.
@@
This principle is operationalized through a suite of two complementary algorithms. The first, Group Token Policy Optimization (GTPO), is a novel token-level algorithm that assigns a unique, entropy-weighted reward to every token, achieving the first true, fine-grained, per-token credit assignment within the efficient GRPO framework. Complementing this, Sequence-Level Group Relative Policy Optimization (GRPO-S) is a lightweight variant that modulates the global reward for an entire sequence based on its average entropy. Together, these methods offer a principled trade-off between granular precision and computational cost, as illustrated in Fig. 1, which conceptually demonstrates how entropy modulation at both token and sequence levels leads to more nuanced credit assignment and improved performance.
@@
  • We formalize coarse-grained credit assignment as a fundamental bottleneck in value-function-free RL, demonstrating its impact on learning efficiency for long-chain reasoning.
@@
  • We provide a theoretical analysis, rooted in variance reduction arguments, to motivate our objective function design and conduct comprehensive experiments on challenging reasoning benchmarks, showing that our methods significantly outperform strong baselines.
@@ 2 DYNAMIC ENTROPY WEIGHTING FOR POLICY OPTIMIZATION
This section introduces Dynamic Entropy Weighting, a framework that addresses coarse-grained credit assignment by repurposing policy entropy for fine-grained reward shaping, motivated by the statistical limitations of GRPO (§ 2.1). Two algorithms are presented: the token-level Group Token Policy Optimization (GTPO) (§ 2.2) and the sequence-level GRPO-S (§ 2.3), and conclude with an analysis of implementation and convergence (§ 2.4).
Figure 2: A high-level comparison of the reward signaling process. Conventional methods like GRPO/DAPO use a static reward model to assign a uniform reward to an entire sequence. Our framework, encompassing GTPO and GRPO-S, introduces a Dynamic Entropy Weighting module that reshapes this signal into fine-grained rewards at either the token or sequence level before it is used by the policy model.
@@ 2.1 From Coarse-Grained Credit Assignment to Dynamic Entropy Weighting
$\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r_{k}\}_{k=1}^{G})}{\text{std}(\{r_{k}\}_{k=1}^{G})}.$ (1)
The GRPO objective then applies this uniform advantage estimate to every token in the sequence within a PPO-style clipped loss function (Shao et al., 2024). While simple and effective, this uniform application of the advantage is the central mechanism of GRPO, and it also represents the core limitation that motivates our work: coarse-grained credit assignment. This approach is not only conceptually imprecise but, as will now be demonstrated, also statistically suboptimal.
@@
$\textbf{Sequence-level mean:}\hat{R}_{1}=\frac{1}{G}\sum_{i=1}^{G}r_{i}.\qquad\textbf{Token-level mean:}\hat{R}_{2}=\frac{\sum_{i=1}^{G}|o_{i}|r_{i}}{\sum_{i=1}^{G}|o_{i}|}.$
While a token-level reward baseline provably reduces variance for more stable gradients ($\text{Var}(\hat{R}_{2})\leq\text{Var}(\hat{R}_{1})$, see Appendix B.1 for a formal proof), this statistical license is insufficient. To fully exploit this granularity, a principled reshaping of the reward signal itself is essential. This is achieved through dynamic entropy weighting, which, as illustrated in Fig. 2, focuses the policy gradient on critical decision points to create a far more instructive and fine-grained learning signal than conventional methods.
Solution: The Dynamic Entropy Weighting Framework. Our framework is built upon the hypothesis that high-entropy moments within a reasoning sequence are not noise but signatures of pivotal junctures. Policy entropy, traditionally a measure of uncertainty, is repurposed as a heuristic for ”cognitive effort”, transforming the sparse, binary reward into a dense, fine-grained learning signal. The framework partitions sequences by their terminal reward into successful ($O^{+}$) and unsuccessful ($O^{-}$) sets, enabling a dual strategy for credit assignment. High-entropy tokens in successful sequences ($o_{i}\in O^{+}$) receive a reward bonus to reinforce valuable exploration. Conversely, low-entropy tokens in unsuccessful sequences ($o_{j}\in O^{-}$) are assigned larger penalties to discourage confident but incorrect reasoning. This precise modulation focuses the policy gradient on the most informative steps of the reasoning process.
@@ 2.2 Group Token Policy Optimization (GTPO)
Group Token Policy Optimization (GTPO) is the most direct and granular implementation of our framework. It introduces a fine-grained, entropy-weighted credit assignment mechanism that operates at the individual token level.
@@
$\tilde{r}_{i,t}^{+}=\alpha_{1}r_{i}+\alpha_{2}\frac{H_{i,t}}{\sum_{k=1}^{n}H_{k,t}}\cdot d_{t}\quad\text{and}\quad\tilde{r}_{j,t}^{+}=0.$ (2)
This reward is composed of the original binary success signal $r_{i}$ (where $r_{i}=1$) and a dynamic entropy bonus, balanced by hyperparameters $\alpha_{1},\alpha_{2}>0$. The bonus is proportional to the token’s generation entropy, $H_{i,t}=-\sum_{v\in\mathcal{V}}\pi_{\theta_{\text{old}}}(v|q,o_{i,<t})\log\pi_{\theta_{\text{old}}}(v|q,o_{i,<t})$. Crucially, this entropy is normalized across all $n$ successful sequences at timestep $t$, creating a relative signal that rewards valuable exploration—tokens generated with higher uncertainty compared to alternative successful paths (if a sequence $o_{k}$ has length less than $t$, its $H_{k,t}$ is treated as 0). This relative bonus is then scaled by $d_{t}$, the count of successful sequences with length $\geq t$, which dynamically adjusts the reward magnitude to account for the diminishing number of active reasoning paths over time.
For any token $o_{j,t}$ within an unsuccessful sequence $o_{j}\in O^{-}$, the goal is to penalize confident mistakes more heavily. Its reward $\tilde{r}_{j,t}^{-}$ is thus defined using inverse entropy, which assigns a larger penalty to low-entropy (i.e., high-confidence) tokens:
$\tilde{r}_{j,t}^{-}=\alpha_{1}\cdot(-1)+\alpha_{2}\frac{1/H_{j,t}}{\sum_{k=1}^{m}(1/H_{k,t})}\cdot h_{t}\cdot(-1)\quad\text{and}\quad\tilde{r}_{i,t}^{-}=0,$ (3)
where $h_{t}$ is the count of unsuccessful sequences with length $\geq t$. This formulation encourages the model to be uncertain when it is incorrect, promoting exploration away from failure modes. Based on these shaped rewards, separate advantage functions are computed for the positive and negative sets, both normalized over all tokens in the entire batch to ensure a consistent scale:
@@
Here, $\tilde{R}^{+}$ and $\tilde{R}^{-}$ represent the collections of all shaped token rewards across all positive and negative sequences in the batch, respectively.
The GTPO Objective Function. The final objective function for GTPO integrates these components into a unified, token-level PPO-style loss. The expectation is taken over all tokens in the batch, weighted by the reciprocal of the total number of tokens $1/\sum_{k=1}^{G}|o_{k}|$:
$\begin{split}\mathcal{J}_{\text{GTPO}}(\theta)=\mathbb{E}\bigg[\frac{1}{\sum_{k=1}^{G}|o_{k}|}\bigg(&\sum_{i=1}^{n}\sum_{t=1}^{|o_{i}|}\min\big(w_{i,t}(\theta)\tilde{A}_{i,t}^{+},\text{clip}(w_{i,t}(\theta),1-\epsilon,1+\epsilon)\tilde{A}_{i,t}^{+}\big)\\ +&\sum_{j=n+1}^{G}\sum_{t=1}^{|o_{j}|}\min\big(w_{j,t}(\theta)\tilde{A}_{j,t}^{-},\text{clip}(w_{j,t}(\theta),1-\epsilon,1+\epsilon)\tilde{A}_{j,t}^{-}\big)\bigg)\bigg],\end{split}$ (5)

### v6

@@ Abstract
Reinforcement Learning (RL) is pivotal for enhancing Large Language Model (LLM) reasoning, yet mainstream algorithms such as GRPO and DAPO remain constrained by a coarse-grained credit assignment paradigm, where all tokens within the same response receive the identical reward. In this paper, we propose Dynamic Entropy Weighting, systematically define entropy-based weight ratios $\frac{H_{i,t}}{\sum_{k=1}^{n}H_{k,t}}$ and similar variants to redistribute rewards and get fine-grained rewards through two new algorithms: Group Token Policy Optimization (GTPO), which assigns an entropy-weighted reward to each token and synthesizes token-specific advantage function to drive the model toward optimal path, and the analogous algorithm Sequence-Level GRPO (GRPO-S), which extends this design to the sequence level and exhibits superior stability in long Chain-of-Thought (CoT) reasoning tasks. Unlike methods using entropy as mere regularization, GTPO and GRPO-S establish a new state-of-the-art on AIME and MATH 500, outperforming prior entropy-guided baselines and validating our weighting mechanism.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Conceptual illustration of reward assignment. (a) Previous methods assign a uniform reward based on the final outcome. In contrast, our methods use Dynamic Entropy Weighting to refine credit assignment: (b) GTPO rewards high-entropy tokens in correct sequences while suppressing them in incorrect ones, and (c) GRPO-S rewards correct sequences with higher average entropy while penalizing incorrect paths, yielding superior performance on AIME 2024 (left) and AIME 2025 (right) (d).
Large-scale Reinforcement Learning (RL) propels LLMs from pattern recognition to complex reasoning, enabling dynamic, multi-step problem-solving (Guo et al., 2025; Wei et al., 2022; Zhang et al., 2022; Achiam et al., 2023; Kimi et al., 2025; Jin et al., 2025). Standard RL approaches optimize policies by maximizing cumulative rewards, typically relying on a value function to estimate expected future returns. However, due to the computational burden of maintaining such critic models, the field evolves from resource-intensive PPO-based pipelines (Mnih et al., 2015; Amini et al., 2024; Christiano et al., 2017; OpenAI Spinning Up, 2018; Gao et al., 2023) to streamlined, value-function-free paradigms. Exemplified by methods such as GRPO, this approach bypasses explicit critic models by leveraging outcome-based group baselines, thereby reducing the computational overhead for massive models (Shao et al., 2024; Yu et al., 2025; Feng et al., 2025b, a).
Despite the remarkable performance of algorithms like GRPO in mathematical and long Chain-of-Thought (CoT) reasoning tasks, a critical bottleneck remains: coarse-grained credit assignment (Uesato et al., 2022; Wei et al., 2022). Algorithms like GRPO assign a uniform reward to every token in a sequence based solely on the final outcome (Shao et al., 2024). This limitation becomes profound in tasks requiring long-chain reasoning (Cui et al., 2025a; Rafailov et al., 2024). For instance, a sequence with dozens of correct logical steps may receive zero reward for a single final error, penalizing correct and incorrect reasoning alike (Yang et al., 2025; Wang et al., 2025b). Conversely, a sequence that reaches a correct answer through flawed or guessed intermediate steps is fully rewarded (Sutton & Barto, 2018; Zhang et al., 2025).
We identify policy entropy as an intrinsic anchor for fine-grained credit assignment, distinguishing our work from recent approaches that treat it merely as a regularization constraint (Cui et al., 2025b) or a binary selection mask (Wang et al., 2025b). While these methods recognize that high entropy signals pivotal reasoning steps (Cheng et al., 2025), they perform passive filtering and fail to address the bottleneck of coarse-grained rewards. In contrast, our framework repurposes entropy for active reward shaping. Instead of discarding tokens, we dynamically redistribute the reward signal. By reinforcing high-entropy exploration in correct solutions while penalizing confident errors in incorrect ones, we achieve precise token-level supervision without an explicit value network.
@@
The central observation is that moments of high policy entropy within a reasoning sequence are not random noise but strong correlates of pivotal reasoning junctures (Cheng et al., 2025; Cui et al., 2025b). When a model selects between multiple valid mathematical theorems or constructs a complex logical connective, its uncertainty, as measured by entropy, naturally increases. This policy entropy, viewed as a measure of model indecision, can be repurposed as a powerful heuristic for cognitive effort (Haarnoja et al., 2018; Lindsay, 2020). In successful paths, it signals a moment of valuable exploration to be reinforced; in unsuccessful paths, it can help the policy break from incorrect thinking. This principle motivates the proposal of Dynamic Entropy Weighting, a novel framework that reshapes the reward signal to be proportional to token-level or sequence-level entropy, thereby focusing the policy gradient on the most critical decision points.
@@
We operationalize this framework through a suite of two complementary algorithms. The first, Group Token Policy Optimization (GTPO), is a novel token-level algorithm that assigns a unique, entropy-weighted reward to every token, achieving the first true, fine-grained, per-token credit assignment within the efficient GRPO framework. Complementing this, Sequence-Level Group Relative Policy Optimization (GRPO-S) is a lightweight variant that modulates the global reward for an entire sequence based on its average entropy. Together, these methods offer a principled trade-off between micro-level granular precision and macro-level sequence stability, as illustrated in Fig. 1, which conceptually demonstrates how entropy modulation at both token and sequence levels leads to more nuanced credit assignment and improved performance.
@@
  • We demonstrate that GRPO still performs coarse-grained credit assignment, and motivated by this, we fundamentally optimize the GRPO framework.
@@
  • We provide a theoretical analysis of unbiasedness and convergence to motivate our objective functions design and conduct comprehensive experiments on challenging reasoning benchmarks, showing that our methods significantly outperform strong baselines.
@@ 2 Dynamic Entropy Weighting For Policy Optimization
We now present the Dynamic Entropy Weighting framework in detail. Motivated by the statistical limitations of GRPO (§ 2.1), we present the detailed formulations of GTPO (§ 2.2) and GRPO-S (§ 2.3) with a theoretical analysis regarding the unbiasedness and convergence (§ 2.4).
Figure 2: A high-level comparison of the reward signaling process. Methods such as GRPO and DAPO use a static reward model to assign a uniform reward to an entire sequence. Our framework, encompassing GTPO and GRPO-S, introduces a Dynamic Entropy Weighting module that reshapes this signal into fine-grained rewards at either the token or sequence level before it is used by the policy model.
@@ 2.1 From Coarse-Grained Credit Assignment to Dynamic Entropy Weighting
$\hat{A}_{i,t}=\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r_{k}\}_{k=1}^{G})}{\text{std}(\{r_{k}\}_{k=1}^{G})}.$ (1)
The GRPO objective then applies this uniform advantage estimate to every token in the sequence within a PPO-style clipped loss function:
$\begin{aligned} \mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}\bigg[\frac{1}{G}&\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\bigg(w_{i,t}(\theta)\hat{A}_{i,t},\text{clip}(w_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\bigg)\bigg].\end{aligned}$
While simple and effective, this uniform application of the advantage is the central mechanism of GRPO, and it also represents the core limitation that motivates our work: coarse-grained credit assignment. This approach is not only conceptually imprecise but, as will now be demonstrated, also statistically suboptimal.
@@
Sequence-level mean: $\quad\hat{R}_{1}=\frac{1}{G}\sum_{i=1}^{G}r_{i}.$
Token-level mean: $\quad\hat{R}_{2}=\frac{\sum_{i=1}^{G}|o_{i}|r_{i}}{\sum_{i=1}^{G}|o_{i}|}.$
While a token-level reward baseline provably reduces variance for more stable gradients $($ $\text{Var}(\hat{R}_{2})\leq\text{Var}(\hat{R}_{1})$, see Appendix B.1 for a formal proof )). It is easy to see that the design of GRPO essentially estimates reward by Sequence-level mean $\hat{R}_{1}$, but according to the analysis in Appendix B.1 & B.2, we know that Token-level mean $\hat{R}_{2}$ is better. However, this exposes a critical issue: GRPO exhibits a dependency on the length of the response sequence. Specifically, for a correct response, it implies that ”the longer, the better”, which is obviously incorrect. Therefore, in a sense, GRPO still performs coarse-grained credit assignment. The underlying cause of this issue is that the design of the reward signal leads to an advantage function that fails to capture the specificity of each token.
Solution: The Dynamic Entropy Weighting Framework. To fully exploit this granularity, a principled reshaping of the reward signal itself is essential. This is achieved through dynamic entropy weighting, which, as illustrated in Fig. 2, focuses the policy gradient on critical decision points to create a far more instructive and fine-grained learning signal than uniform credit assignment approaches.
The framework partitions sequences (in $O$, where $O$ is the set consists of all response sequences) by their terminal reward into successful ($O^{+}$) and unsuccessful ($O^{-}$) sets, enabling a dual strategy for credit assignment. High-entropy tokens in successful sequence ($o_{i}\in O^{+}$) receive a reward bonus to reinforce valuable exploration. Conversely, low-entropy tokens in unsuccessful sequence ($o_{j}\in O^{-}$) are assigned relatively larger penalties to discourage confident but incorrect reasoning. This precise modulation focuses the policy gradient on the most informative steps.
@@ 2.2 Group Token Policy Optimization (GTPO)
Group Token Policy Optimization (GTPO) is the most direct and granular implementation of our framework. It introduces a fine-grained, entropy-weighted credit assignment mechanism that operates at the token level.
@@
$\tilde{r}_{i,t}^{+}=\alpha_{1}r_{i}+\alpha_{2}\frac{H_{i,t}}{\sum_{k=1}^{n}H_{k,t}}\cdot d_{t},\quad\quad\tilde{r}_{j,t}^{+}=0.$ (3)
This reward is composed of the original binary success signal $r_{i}$ (where $r_{i}=1$) and a dynamic entropy bonus, balanced by hyperparameters $\alpha_{1},\alpha_{2}>0$. The bonus is proportional to the token’s generation entropy
$H_{i,t}=-\sum_{v\in\mathcal{V}}\pi_{\theta_{\text{old}}}(v|q,o_{i,<t})\log\pi_{\theta_{\text{old}}}(v|q,o_{i,<t}).$ (4)
Crucially, this entropy is normalized across all $n$ successful sequences at timestep $t$, creating a relative signal that rewards valuable exploration, specifically targeting tokens generated with higher uncertainty compared to alternative successful paths (if a sequence $o_{i}$ has length less than $t$, its $H_{k,t}$ is treated as 0). This relative bonus is then scaled by $d_{t}$, the count of successful sequences with length $\geq t$, which dynamically adjusts the reward magnitude to account for the diminishing number of active reasoning paths over time.
For any token $o_{j,t}$ within an unsuccessful sequence $o_{j}\in O^{-}$, the goal is to penalize confident mistakes more heavily. The reward $\tilde{r}_{j,t}^{-}$ is thus defined using inverse entropy to assign relatively larger penalties to low-entropy (i.e., high-confidence) tokens:
$\begin{split}\tilde{r}_{j,t}^{-}&=\alpha_{1}\cdot(-1)+\alpha_{2}\frac{1/H_{j,t}}{\sum_{k=1}^{m}(1/H_{n+k,t})}\cdot h_{t}\cdot(-1),\\ \tilde{r}_{i,t}^{-}&=0.\end{split}$ (5)
where $h_{t}$ is the count of unsuccessful sequences with length $\geq t$ (if a sequence $o_{j}$ has length less than $t$, its $1/{H}_{j,t}$ is treated as 0). This formulation encourages the model to be uncertain when it is incorrect, promoting exploration away from failure modes. Based on these shaped rewards, separate advantage functions are computed for the positive and negative sets, both normalized over all tokens in the entire batch to ensure a consistent scale:
@@
Here, $\tilde{R}^{+}$ and $\tilde{R}^{-}$ represent the collections of all shaped token rewards across all positive and negative sequences in the group, respectively.
The GTPO Objective Function. The final objective function for GTPO integrates these components into a unified, token-level PPO-style loss. The expectation is taken over all tokens in the batch, weighted by the reciprocal of the total number of tokens, $1/\sum_{k=1}^{G}|o_{k}|$:
$\begin{aligned} \mathcal{J}_{\text{GTPO}}(\theta)=\mathbb{E}\bigg[\frac{1}{\sum_{k=1}^{G}|o_{k}|}\bigg(&\sum_{i=1}^{n}\sum_{t=1}^{|o_{i}|}\min\big(w_{i,t}(\theta)\tilde{A}_{i,t}^{+},\text{clip}(w_{i,t}(\theta),1-\epsilon,1+\epsilon)\tilde{A}_{i,t}^{+}\big)\\ +&\sum_{j=n+1}^{G}\sum_{t=1}^{|o_{j}|}\min\big(w_{j,t}(\theta)\tilde{A}_{j,t}^{-},\text{clip}(w_{j,t}(\theta),1-\epsilon,1+\epsilon)\tilde{A}_{j,t}^{-}\big)\bigg)\bigg],\end{aligned}$